Steve traveled 200 miles at a certain speed. Had he gone 10 mph faster, the trip would have been 1 hour less. Find the speed of his vehicle.
Answer by jorel1380(3719)

200/x=200/(x+10) +1
200(x+10)=200x+x(x+10)
200x+2000=200x+x2+10x
x2+10x-2000=0
(x+ 50)(x-40)=0
x=40 or -50
Throwing out the negative answer, we get Steve's speed to be 40 mph..